OpenStudy (anonymous):

Can someone take me through this problem? -Create an exponential growth function, f(x), to model a population of frogs that is growing every year. Identify the principal amount, the growth rate, and the appropriate domain and range for your function. Explain how these key features would affect the graph of f(x).

OpenStudy (anonymous):

I came up with this function: f(x)=100*(0.05)^5 100 = the initial frog population 0.05 = that they grow 5% per year ^5 = 5 years Did I do this correctly?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

For 5% growth it should be f(x)=100*(\(\color{red}{1}\).05)^5

ganeshie8 (ganeshie8):

Also, replace 5 by x because the variable "x" represents the number of years : f(x)=100*(\(\color{red}{1}\).05)^\(\color{red}{x}\)

ganeshie8 (ganeshie8):

Other than that, your work looks fine! good job!

OpenStudy (anonymous):

Oh because we don't know how many years

OpenStudy (anonymous):

Alright how would I create the domain and range?

ganeshie8 (ganeshie8):

Kindof, it is because we can think of the function "f(x)" as a formula to which you feed in the number of years "x", and it spits out the population of frogs in that year.